# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v4

@@ 4.9.2 FCC Mobility Law
All dislocation segments are constrained to the $\{111\}$ slip planes in FCC, and their respective drag tensor can be expressed as [90]
@@ 4.9.3 Incorporating finite boundary conditions
To evaluate the corrective force term $\boldsymbol{\hat{F}}^{k}$ in Equation (10), the superposition principle is adopted. As described elsewhere [91, 92], the total stress strain and displacement fields are expressed as
@@
respectively. The infinite-body fields are denoted as $\tilde{()}$ and the finite-element correction fields as $\hat{()}$. According to the following procedure [93], the image stress field may be evaluated. First, the elastic stress field due to dislocations in an infinite body, $\tilde{\sigma}$, is obtained. The tractions $\boldsymbol{\tilde{T}}=\boldsymbol{\tilde{\sigma}}\cdot n$ on the traction boundaries due to this stress are then calculated, and subtracted from the desired boundary conditions $\boldsymbol{T}$. These modified boundary values, $\boldsymbol{\hat{T}}=\boldsymbol{T}-\boldsymbol{\tilde{T}}$, in addition to the displacement conditions $\boldsymbol{U}$ on the displacement boundaries are used in an elastic finite element simulation to determine the corrective fields. Finally, the corrective stress field, $\boldsymbol{\hat{\sigma}}$, is used to evaluate the corrective nodal force, $\boldsymbol{\hat{f}}$.
@@ 4.10 Dislocation strain field modeling
It was assumed that there were no externally applied stresses. The $\varepsilon_{111,\ \mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,\ t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [94, 95], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
@@ Acknowledgements
F.H. thanks Thomas D. Swinburne and Max Boleininger for insightful comments. D.Y., G.H., and F.H. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 714697). K.S. acknowledges funding from the General Sir John Monash Foundation. A.M. acknowledges funding from EDF (Électricité de France). Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. This research used the Materials Synthesis and Characterization Facility of the Center for Functional Nanomaterials (CFN), which is a U.S. Department of Energy Office of Science User Facility, at Brookhaven National Laboratory under Contract No. DE-SC0012704. Experiments were performed at the Advanced Photon Source, a US Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357. Work performed at UCL was supported by EPSRC. The authors acknowledge the use of the Advanced Research Computing (ARC) facility at the University of Oxford [96].
@@
The data that support the findings of this study are openly available in Zenodo [97].

### v5

@@ 4.9.2 FCC Mobility Law
All dislocation segments are constrained to the $\{111\}$ slip planes in FCC, and their respective drag tensor can be expressed as [Bulatov:NA:2006]
@@ 4.9.3 Incorporating finite boundary conditions
To evaluate the corrective force term $\boldsymbol{\hat{F}}^{k}$ in Equation (10), the superposition principle is adopted. As described elsewhere [Giessen:MSMSE:1995, Weygand:MSMSE:2002], the total stress strain and displacement fields are expressed as
@@
respectively. The infinite-body fields are denoted as $\tilde{()}$ and the finite-element correction fields as $\hat{()}$. According to the following procedure [El-Awady:JotMaPoS:2008], the image stress field may be evaluated. First, the elastic stress field due to dislocations in an infinite body, $\tilde{\sigma}$, is obtained. The tractions $\boldsymbol{\tilde{T}}=\boldsymbol{\tilde{\sigma}}\cdot n$ on the traction boundaries due to this stress are then calculated, and subtracted from the desired boundary conditions $\boldsymbol{T}$. These modified boundary values, $\boldsymbol{\hat{T}}=\boldsymbol{T}-\boldsymbol{\tilde{T}}$, in addition to the displacement conditions $\boldsymbol{U}$ on the displacement boundaries are used in an elastic finite element simulation to determine the corrective fields. Finally, the corrective stress field, $\boldsymbol{\hat{\sigma}}$, is used to evaluate the corrective nodal force, $\boldsymbol{\hat{f}}$.
@@ 4.10 Dislocation strain field modeling
It was assumed that there were no externally applied stresses. The $\varepsilon_{111,\ \mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,\ t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [Barnett1985, Barnett2007], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
@@ Acknowledgements
F.H. thanks Thomas D. Swinburne and Max Boleininger for insightful comments. D.Y., G.H., and F.H. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 714697). K.S. acknowledges funding from the General Sir John Monash Foundation. A.M. acknowledges funding from EDF (Électricité de France). Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. This research used the Materials Synthesis and Characterization Facility of the Center for Functional Nanomaterials (CFN), which is a U.S. Department of Energy Office of Science User Facility, at Brookhaven National Laboratory under Contract No. DE-SC0012704. Experiments were performed at the Advanced Photon Source, a US Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357. Work performed at UCL was supported by EPSRC. The authors acknowledge the use of the Advanced Research Computing (ARC) facility at the University of Oxford [Richards2015].
@@
The data that support the findings of this study are openly available in Zenodo [zenodo_ref].